### v6

@@ Abstract
We study the bias of classical quantile regression and instrumental variable quantile regression estimators. While being asymptotically first-order unbiased, these estimators can have non-negligible second-order biases. We derive a higher-order stochastic expansion of these estimators using empirical process theory. Based on this expansion, we derive an explicit formula for the second-order bias and propose a feasible bias correction procedure that uses finite-difference estimators of the bias components. The proposed bias correction method performs well in simulations. We provide an empirical illustration using Engel’s classical data on household expenditure.
@@
Keywords: instrumental variables, higher-order stochastic expansion, Bahadur-Kiefer expansion, finite-difference estimators, mixed integer linear programming (MILP).
@@ 1 Introduction
To study the biases, we start by deriving a higher-order stochastic expansion of the classical QR and exact IVQR estimators. [^2] We define exact IVQR estimators as estimators that exactly minimize a norm of the sample moment conditions. Such estimators can be obtained using mixed-integer programming (MIP) methods (e.g., Chen and Lee, 2018; Zhu, 2019). See Appendix D. Such an expansion is needed because the higher-order terms contribute nonzero biases while the first-order term does not. We derive explicit expressions and uniform (in the quantile level) rates for the components in the expansion, building on the empirical process arguments of Ota et al. (2019). This expansion can be thought of as a refined Bahadur-Kiefer (BK) representation of the estimator that decomposes the nonlinear component into terms up to the order $O_{p}\left(n^{-1}\right)$ and a $O_{p}\left(n^{-5/4}\sqrt{\log n}\right)$ remainder. [^3] In Appendix E, we also derive a uniform BK representation for generic IVQR estimators after a feasible 1-step Newton correction.
@@
Using the stochastic expansion, we study the bias of QR and exact IVQR estimators. We derive a bias formula based on the leading terms up to order $O_{p}\left(n^{-1}\right)$ in the expansion, which we refer to as the second-order (asymptotic) bias. This approach of focusing on the moments of the leading terms in the stochastic expansions is standard in the literature. [^4] See, for example, Nagar (1959); Newey and Smith (2004); Kato et al. (2012); Galvao and Kato (2016); Kaplan and Sun (2017); Hahn et al. (2023), among others. The second-order bias formula provides an approximation of the actual bias that yields a feasible correction. Our results explicitly account for the $O\left(n^{-1}\right)$ (up to a logarithmic factor) bias due to nonzero sample moments at the estimator. Our proof strategy is different from the generalized function heuristic used in the existing literature (Phillips, 1991; Lee et al., 2017, 2018), which does not account for all the leading terms up to $O\left(n^{-1}\right)$ in the bias. [^5] See Section 3.2 and Appendix C for further discussion and examples. The missing terms can be important bias contributors in practice, as we document in the empirical application in Section 5.
@@
A feasible bias correction procedure then follows from the second-order bias formula. We propose finite-difference estimators of all the components in the formula. These estimators admit higher-order expansions that allow us to select bandwidth rates. [^6] In particular, our finite-difference estimator for the Jacobian coincides with Powell (1986)’s classical estimator, and the bandwidth rate we derive coincides with the AMSE optimal bandwidth choice in Kato (2012). We show that the resulting bias-corrected estimator has zero second-order bias.
@@
We evaluate the performance of our bias correction procedure in a Monte Carlo simulation study. The simulations show that the theoretical (infeasible) bias formula describes well the second-order bias of classical QR and exact IVQR. We find that the proposed feasible bias correction can effectively reduce the bias in many cases, even in samples as small as $n=50$. However, strong asymmetry and, especially, heavy tails in the conditional distribution of the outcome may reduce the effectiveness of bias correction. The impact of the feasible bias correction on the root MSE (RMSE) is rather small and ambiguous.
@@
We illustrate the bias correction approach by revisiting the relationship between food expenditure and income based on the original Engel (1857) data (e.g., Koenker and Bassett, 1982; Koenker and Hallock, 2001). Our results highlight the importance of bias correction in empirical applications with small sample sizes. Specifically, we find that the second-order bias of classical QR is non-negligible, especially at the upper tail, leading QR to underestimate the effect heterogeneity across quantiles.
@@ 2 Model and estimators
As noted by Chernozhukov and Hansen (2006, p.502), compactness of the parameter space $\Theta$ “is not restrictive in micro-econometric applications.” Throughout the paper, we use the short notation $G$ for $G(\theta_{\tau})$ whenever it does not lead to ambiguity.
@@
Assumption 3 guarantees the existence of all relevant moments of the terms involving $W$ and $Z$ in the higher-order derivatives of the moment conditions and the bias correction (e.g., $\mathbb{E}Z_{\ell}W_{j}W_{q}W_{r}$). The power parameter $\gamma$ (as we show below) determines the upper bound on the rate at which the norm of the sample moment functions converge to zero — higher $\gamma$ implies faster convergence to zero.
@@ 3 Asymptotic theory for bias correction
To derive a bias correction procedure, we follow the approach of Nagar (1959) and focus on the bias of the leading terms in an asymptotic stochastic expansion of the estimator.
@@ 3.1 Stochastic expansion of quantile regression estimators
To state the result, we introduce some additional notation. For $\theta\in\Theta$, define the auxiliary functions $g^{\circ}(\theta)\triangleq\mathbb{E}1\{Y\leq W^{\prime}\theta\}Z$, $B^{\circ}_{n}(\theta)\triangleq\sqrt{n}(\mathbb{E}_{n}1\{Y\leq W^{\prime}\theta\}Z-g^{\circ}(\theta))$, and $B_{n}(\theta)\triangleq B^{\circ}_{n}(\theta)-B^{\circ}_{n}(\theta_{\tau})$. [^7] Processes $B_{n}(\theta)$ and $B^{\circ}_{n}(\theta)$ take values in the space $\ell^{\infty}(\Theta)$ of bounded functions on $\Theta$. Also, denote the Hessian of the $j$-th moment function $g_{j}$ by $\partial_{\theta}G_{j}(\theta)\triangleq\partial_{\theta}\partial_{\theta}g_{j}(\theta)$. Finally, for any $x\in\mathbb{R}^{k}$, denote by $x^{\prime}\partial_{\theta}G(\theta)x$ the vector with components $x^{\prime}\partial_{\theta}G_{j}(\theta)x$, $j=1,\dots,k$.
@@
[…]
@@ 3.2 Bias formula for exact estimators
The second component, $n^{-1}G^{-1}\kappa_{\tau}$, appears because of the discontinuity in the sample moment functions. The term $\kappa_{\tau}$ reflects the dependence between the sample moments and the linear influence of a single observation on $\hat{\theta}_{\tau}$. Notice that the first two components combined correspond to the second order bias of the terms $G^{-1}(\theta_{\tau})(\hat{g}_{\tau}(\hat{\theta}_{\tau})-n^{-1/2}B_{n}(\hat{\theta}_{\tau}))$ in Theorem 1.
The last component, $(2n)^{-1}G^{-1}Q^{\prime}vec(\Omega)$, stems from the non-uniformity of the conditional distribution of $Y$ given $(W,Z)$. This term correspond to the term $G^{-1}(\theta_{\tau})(\hat{\xi}_{\tau}-\theta_{\tau})^{\prime}\partial_{\theta}G(\theta_{\tau})(\hat{\xi}_{\tau}-\theta_{\tau})/2$ in Theorem 1. Similar terms are typically present in most nonlinear estimators with nonzero Hessian of the score function (see, for example, Rilstone et al., 1996).
@@
Figure 1: Exact (circles) and second-order (crosses) biases, scaled by $n$, as functions of quantile level $\tau$ for $\hat{\theta}_{\tau}=Y_{(\lfloor\tau n\rfloor)}$, where $Y\sim\text{Uniform}(0,1)$, $n=10$.
@@ 3.3 Feasible bias correction
Suppose that Assumptions 14 hold. Consider $\hat{\theta}_{\tau}=\hat{\theta}_{\tau,p}$ obtained from program (3) for some $p\in[1,\infty]$ or $\hat{\theta}_{\tau}=\hat{\theta}_{\tau,QR}$. Then the bias correction eliminates the second-order bias, $\operatorname{Bias}\left(\hat{\theta}_{bc}\right)=0.$
@@ 3.4 Finite difference estimators of bias components
Suppose that Assumptions 13 hold. Then the nearly remainder-optimal bandwidth rates are $h_{1,n}\propto n^{-1/5}$, $h_{2,n}\propto n^{-1/7}$, $h_{3,n}\propto n^{-2/15}$, and, under these bandwidth rates,
@@
$\hat{\kappa}_{\tau}$ $=\kappa_{\tau}+O_{p}\left(\frac{\sqrt{\log n}}{n^{4/15}}\right),$
@@
Proposition 1 in Kato (2012) shows that the remainder rate for $\hat{G}$ in Lemma 1, $h_{1,n}\propto n^{-1/5}$, is the AMSE-optimal rate.888 We conjecture that analogous AMSE-optimality results could be established for estimators $\widehat{({\partial_{\theta}G}_{i,j})}_{\ell}$ and $\hat{Q}_{j}$, but leave this extension for future work. To implement the finite difference estimators in practice, one needs to choose constants in addition to the bandwidth rates. We discuss the choice of these constants in Sections 4 and 5.
@@ 4 Simulation evidence
In this section, we evaluate the performance of our feasible bias correction procedure in a Monte Carlo simulation study. We consider data-generating processes (DGPs) inspired by the simulations in Andrews and Mikusheva (2016). The outcome is generated according to the following location-scale model
@@
where $W_{i}=\Phi(\tilde{W}_{i})$, $Z_{i}=\Phi(\tilde{Z}_{i})$, $U_{i}=F_{U}^{-1}\left(\Phi(\tilde{U}_{i})\right)$, $(\tilde{W}_{i},\tilde{Z}_{i},\tilde{U}_{i})\sim N(0,\Sigma)$, $\Sigma_{11}=\Sigma_{22}=\Sigma_{33}=1$, $\Sigma_{23}=0$, and $\Phi$ is the standard normal CDF. Hence, in all the designs, both the regressors and the instruments are $\operatorname{Uniform}[0,1]$. We consider four DGPs that differ with respect to the error distribution $F_{U}$ and whether or not $W$ is exogenous.
@@
An important practical issue is the choice of the bandwidths $h_{1,n}$, $h_{2,n}$, and $h_{3,n}$. We choose the nearly remainder-optimal rates for the bandwidths from Lemma 1: $h_{1,n}=A_{G}n^{-1/5}$, $h_{2,n}=A_{Q}n^{-1/7}$, $h_{3,n}=A_{\kappa}n^{-2/15}$, where $(A_{G},A_{Q},A_{\kappa})$ are constants that are independent of the sample size. We report results for both our baseline choice $(A_{G},A_{Q},A_{\kappa})=(1,1,1)$ and the choice of $(A_{G},A_{Q},A_{\kappa})$ that minimizes the $\ell_{2}$-distance between the residual biases after infeasible and feasible correction for the intercept and the slope parameter over the grid
$\mathcal{A}\triangleq\{(1,1,1),(1.5,1,1),(0.5,1,1),(1,1.5,1),(1,0.5,1),(1,1,1.5),(1,1,0.5)\}.$ (10)
Figure 6 in Appendix F shows how the results change if we increase/decrease each element of $(A_{G},A_{Q},A_{\kappa})$, one at a time.
We focus on the performance of bias correction for $\tau\in\{0.25,0.5,0.75\}$. To evaluate the precision of the Monte Carlo integration, we compute the Monte Carlo standard error (MCSE). We report $\Phi^{-1}(1-0.05/12)\times$ MCSE to account for the joint testing of six hypotheses (two hypotheses per each of three levels of $\tau$). For comparison, we also display the infeasible bias correction based on the true $G$, $Q$, and $\kappa$.
Figure 2 shows the results for the exogenous DGP1 with $n\in\{50,100,200\}$ based on classical QR of $Y$ on $W$, implemented via linear programming (see Appendix D). Notice that the scaled bias of the QR estimator, $n(\mathbb{E}\hat{\theta}_{\tau,\text{QR}}-\theta_{\tau})$, (blue dots) has the same range across the different sample sizes. This illustrates that we are estimating $O(n^{-1})$ terms in the asymptotic bias expansion, which should be approximately constant after scaling by $n$. One can see that the infeasible bias correction (gold dashed lines) reduces the bias at most quantile levels. The correction does not result in zero bias due to the finite precision of the simulation (the sample moment term does not have an explicit bias formula) and the presence of the higher-order bias terms. Notice that the MCSE grows with the sample size for a fixed number of simulations due to the rescaling, which helps explain the seemingly better performance for $n=50$.
We find that the feasible bias correction (gold squares and crosses) typically reduces the bias at $\tau\in\{0.25,0.75\}$, where the bias of classical QR is largest, and can get very close to the infeasible bias correction. At $\tau=0.5$, the bias of classical QR is small so that the feasible bias correction has a negligible impact. The feasible bias correction with the baseline bandwidth choice $(1,1,1)$ results in comparable deviations from the infeasible one across the different sample sizes. This is consistent with Lemma 1 since, otherwise, different sample sizes would require different constants $A_{G},A_{Q},A_{\kappa}$.
In Figure 3, we study the role of the error distribution $F_{U}$ and endogeneity. The infeasible bias correction reduces the bias across all the DGPs. Panels (a) and (b) document two cases that present challenges for the feasible bias correction procedure in small samples: strong asymmetry in DGP2, which may affect the performance at the median, and, especially, heavy tails in DGP3, which may affect the performance at tail quantiles. Panel (c) demonstrates that the bias correction performs well when combined with IVQR implemented using the MILP formulation in Appendix D. While the scaled second-order bias of IVQR can be substantial, especially for the slope parameter, the feasible bias correction effectively reduces that bias across all quantile levels considered.
We investigate the impact of bias correction on the RMSE of the estimators in Figure 4. We report the results for QR based on the exogenous DGP1 and IVQR based on the endogenous DGP4 with $n=100$. Overall, the impact of bias correction on the RMSE is small. While it slightly increases the RMSE for DGP1, it sometimes decreases the RMSE for DGP4. The alternative measure of dispersion, mean absolute deviation, shows similar performance. See Figure 7 in the Appendix.
Finally, we study the impact of the bias correction on the coverage probability of the standard confidence intervals. We use the same asymptotic variance estimator of $\hat{\theta}_{\tau}$, $\hat{G}^{-1}\hat{\Omega}(\hat{G}^{\prime})^{-1}$, as in Section 3.3, for both original and bias-corrected estimator. As a result, by construction, the bias correction does not impact the length of the confidence intervals. Figures 8 and 9 in the Appendix show coverage before and after the bias correction. The bias correction brings the empirical coverage closer to the nominal coverage in many but not all cases. This impact is small for all the DGPs, except for extreme quantiles in DGP3 with heavy tails where the feasible bias correction would require larger sample sizes. (The infeasible formula still results in bias improvement in that case, see Panel (b) in Figure 3.)

### v7

@@ Abstract
We study the bias of classical quantile regression and instrumental variable quantile regression estimators. While being asymptotically first-order unbiased, these estimators can have non-negligible second-order biases. We derive a higher-order stochastic expansion of these estimators using empirical process theory. Based on this expansion, we derive an explicit formula for the second-order bias and propose a feasible bias correction procedure that uses finite-difference estimators of the bias components. The proposed bias correction method performs well in simulations. We provide an empirical illustration using Engel’s classical data on household food expenditure.
@@
Keywords: instrumental variables, higher-order stochastic expansion, Bahadur-Kiefer expansion, finite-difference estimators, mixed integer linear programming (MILP), Engel curve.
@@ 1 Introduction
To study the biases, we start by deriving a higher-order stochastic expansion of the classical QR and exact IVQR estimators. [^2] We define exact IVQR estimators as estimators that exactly minimize a norm of the sample moment conditions. Such estimators can be obtained using mixed-integer programming (MIP) methods (e.g., Chen and Lee, 2018; Zhu, 2019). See Appendix D. Such an expansion is needed because the higher-order terms contribute nonzero biases while the first-order term does not. We derive explicit expressions and uniform (in the quantile level) rates for the components in the expansion, building on the empirical process arguments of Ota et al. (2019). This expansion can be thought of as a refined Bahadur-Kiefer (BK) representation of the estimator that decomposes the nonlinear component into terms up to the order $O_{p}\left(n^{-1}\right)$ and a $O_{p}\left(n^{-5/4}\sqrt{\log n}\right)$ remainder. We also derive a uniform BK representation for generic IVQR estimators after a feasible 1-step Newton correction (see Appendix E).
@@
Using the stochastic expansion, we study the bias of QR and exact IVQR estimators. We derive a bias formula based on the leading terms up to order $O_{p}\left(n^{-1}\right)$ in the expansion, which we refer to as the second-order (asymptotic) bias. This approach of focusing on the moments of the leading terms in the stochastic expansions is standard in the literature. [^3] See, for example, Nagar (1959); Newey and Smith (2004); Kato et al. (2012); Galvao and Kato (2016); Kaplan and Sun (2017); Hahn et al. (2023), among others. The second-order bias formula provides an approximation of the actual bias that yields a feasible correction. Our results explicitly account for the $O\left(n^{-1}\right)$ (up to a logarithmic factor) bias due to nonzero sample moments at the estimator. Our proof strategy is different from the generalized function heuristic used in the existing literature (Phillips, 1991; Lee et al., 2017, 2018), which does not account for all the leading terms up to $O\left(n^{-1}\right)$ in the bias (see Section 3.2 and Appendix C for further discussion and examples). The missing terms can be important bias contributors in practice, as we document in the empirical application in Section 5.
@@
A feasible bias correction procedure then follows from the second-order bias formula. We propose finite-difference estimators of all the components in the formula. These estimators admit higher-order expansions that allow us to select bandwidth rates. In particular, our finite-difference estimator for the Jacobian coincides with Powell (1986)’s classical estimator, and the bandwidth rate we derive coincides with the AMSE optimal bandwidth choice in Kato (2012). We show that the resulting (analytically) bias-corrected estimator has zero second-order bias. This result is in contrast to the commonly-used bootstrap bias correction approaches (Horowitz, 2001), which may not capture the higher-order bias terms of quantile estimators (see Knight, 2003).
@@
We evaluate the performance of our bias correction procedure in a Monte Carlo simulation study. The simulations show that the theoretical (infeasible) bias formula describes well the second-order bias of classical QR and exact IVQR. We find that the proposed feasible bias correction can effectively reduce the bias in many cases. The gains from bias correction are particularly prominent in settings with large bias such as for the IVQR estimators under endogeneity. The impact of the bias correction on the root MSE (RMSE) is rather small and ambiguous.
@@
We illustrate the bias correction approach by revisiting the relationship between food expenditure and income based on the original Engel (1857) data (e.g., Koenker and Bassett, 1982; Koenker and Hallock, 2001). [^4] The QR approach was used more recently to analyze changes in heterogeneity in the income elasticity over time of various consumer spending categories in Taylor and Houthakker (2009). Our results highlight the importance of bias correction in empirical applications with small sample sizes. Specifically, we find that the second-order bias of classical QR is empirically relevant: it can be larger than 50% of the standard error, which is substantial given that this is the second-order bias.
@@ 2 Model and estimators
As noted by Chernozhukov and Hansen (2006), “compactness [of the parameter space $\Theta$] is not restrictive in micro-econometric applications” (p.502). Throughout the paper, we use the short notation $G$ for $G(\theta_{\tau})$ whenever it does not lead to ambiguity.
@@
Assumption 3 guarantees the existence of all relevant moments of the terms involving $W$ and $Z$ in the higher-order derivatives of the moment conditions and the bias correction (e.g., $\mathbb{E}Z_{\ell}W_{j}W_{q}W_{r}$). The power parameter $\gamma$ (as we show below) determines the upper bound on the rate at which the norm of the sample moment functions converges to zero — higher $\gamma$ implies faster convergence to zero.
@@ 3 Asymptotic theory for bias correction
To derive a bias correction procedure, we follow the approach of Nagar (1959) and focus on the bias of the leading terms in the asymptotic stochastic expansion of the estimator.
@@ 3.1 Stochastic expansion of quantile regression estimators
To state the result, we introduce some additional notation. For $\theta\in\Theta$, define the auxiliary functions $g^{\circ}(\theta)\triangleq\mathbb{E}1\{Y\leq W^{\prime}\theta\}Z$, $B^{\circ}_{n}(\theta)\triangleq\sqrt{n}(\mathbb{E}_{n}1\{Y\leq W^{\prime}\theta\}Z-g^{\circ}(\theta))$, and $B_{n}(\theta)\triangleq B^{\circ}_{n}(\theta)-B^{\circ}_{n}(\theta_{\tau})$. [^5] The processes $B_{n}(\theta)$ and $B^{\circ}_{n}(\theta)$ take values in the space $\ell^{\infty}(\Theta)$ of bounded functions on $\Theta$. Also, denote the Hessian of the $j$-th moment function $g_{j}$ by $\partial_{\theta}G_{j}(\theta)\triangleq\partial_{\theta}\partial_{\theta}g_{j}(\theta)$. Finally, for any $x\in\mathbb{R}^{k}$, denote by $x^{\prime}\partial_{\theta}G(\theta)x$ the vector with components $x^{\prime}\partial_{\theta}G_{j}(\theta)x$, $j=1,\dots,k$.
@@
$\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{g}_{\tau}(\hat{\theta}_{\tau})\|$ =$=\begin{cases}O_{p}\left(\frac{1}{n}\right),\text{ if }\hat{\theta}_{\tau}=\hat{\theta}_{\tau,\text{QR}},\\ O_{p}\left(\frac{\log n}{n^{1-\frac{2}{\gamma}}}\right),\text{ if }\hat{\theta}_{\tau}=\hat{\theta}_{\tau,p},\end{cases}$ (5)
@@ 3.2 Bias formula for exact estimators
The second component, $n^{-1}G^{-1}\kappa_{\tau}$, appears because of the discontinuity in the sample moment functions. The term $\kappa_{\tau}$ reflects the dependence between the sample moments and the linear influence of a single observation on $\hat{\theta}_{\tau}$. Notice that the first two components combined correspond to the second-order bias of the terms $G^{-1}(\theta_{\tau})(\hat{g}_{\tau}(\hat{\theta}_{\tau})-n^{-1/2}B_{n}(\hat{\theta}_{\tau}))$ in Theorem 1.
The last component, $(2n)^{-1}G^{-1}Q^{\prime}vec(\Omega)$, stems from the non-uniformity of the conditional distribution of $Y$ given $(W,Z)$. This term corresponds to the term $G^{-1}(\theta_{\tau})(\hat{\xi}_{\tau}-\theta_{\tau})^{\prime}\partial_{\theta}G(\theta_{\tau})(\hat{\xi}_{\tau}-\theta_{\tau})/2$ in Theorem 1. Similar terms are typically present in most nonlinear estimators with nonzero Hessian of the score function (see, for example, Rilstone et al., 1996).
@@
Figure 1: Comparison to exact formula in univariate case
Notes: Exact (circles) and second-order (crosses) biases, scaled by $n$, as functions of quantile level $\tau$ for $\hat{\theta}_{\tau}=Y_{(\lfloor\tau n\rfloor)}$, where $Y\sim\text{Uniform}(0,1)$, $n=10$.
@@ 3.3 Feasible bias correction
Suppose that Assumptions 14 hold. Consider $\hat{\theta}_{\tau}=\hat{\theta}_{\tau,p}$ obtained from program (3) for some $p\in[1,\infty]$ or $\hat{\theta}_{\tau}=\hat{\theta}_{\tau,QR}$. Then the feasible bias correction eliminates the second-order bias, $\operatorname{Bias}\left(\hat{\theta}_{bc}\right)=0.$
@@ 3.4 Finite difference estimators of bias components
Suppose that Assumptions 13 hold. Then the nearly remainder-optimal bandwidth rates are $h_{1,n}\propto n^{-1/5}$, $h_{2,n}\propto n^{-1/7}$, $h_{3,n}\propto n^{-1/5}$, and, under these bandwidth rates,
@@
$\hat{\kappa}_{\tau}$ $=\kappa_{\tau}+O_{p}\left(\frac{\sqrt{\log n}}{n^{2/5}}\right),$
@@
Proposition 1 in Kato (2012) shows that the remainder rate for $\hat{G}$ in Lemma 1, $h_{1,n}\propto n^{-1/5}$, is the AMSE-optimal rate.666 We conjecture that analogous AMSE-optimality results could be established for estimators $\widehat{({\partial_{\theta}G}_{i,j})}_{\ell}$ and $\hat{Q}_{j}$, but leave this extension for future work.
To implement the finite difference estimators in practice, one needs to choose constants in addition to the bandwidth rates. A natural approach would be to use non-parametric methods such as cross-validation to estimate the optimal bandwidth parameter. However, such non-parametric methods typically require large sample sizes (e.g. Simonoff, 1996), which makes them impractical for our purposes. Another common approach, which we take here, is to propose rule-of-thumb choices that account for the scale of the data. To be robust to heavy-tailed distributions, we suggest using a robust measure of dispersion. Specifically, we propose a rule-of-thumb choice in form $h_{1,n}=A_{G}\cdot n^{-1/5}$, $h_{2,n}=A_{Q}\cdot n^{-1/7}$, $h_{3,n}=A_{\kappa}\cdot n^{-1/5}$, where, for all $j\in\{G,Q,\kappa\}$,
$A_{j}=\tilde{A}_{j}\cdot 1.48\cdot\widehat{\operatorname{MAD}}_{\tau},$
and $\widehat{\operatorname{MAD}}_{\tau}$ is the estimated median absolute deviation of the $\tau$-quantile residuals. The constant $1.48$ is chosen because for the normal distribution 1.48 times the median absolute deviation is equal to the standard deviation. [^7] Chernozhukov et al. (2013) suggest using $\operatorname{IQR}/1.35$ as a robust estimate of the scale parameter. For the normal distribution, this choice coincides with $1.48\operatorname{MAD}$. We suggest choosing $\tilde{A}_{G}=\tilde{A}_{\kappa}=2$ and $\tilde{A}_{Q}=1.5$ based on our simulation evidence, as we describe in more detail in Section 4.2.
@@ 4 Simulation evidence
In this section, we evaluate the performance of our feasible bias correction procedure in a Monte Carlo simulation study.
4.1 Simulation design
We consider data-generating processes (DGPs) inspired by the simulations in Andrews and Mikusheva (2016). The outcome is generated according to the following location-scale model
@@ 4.1 Simulation design
where $W_{i}=\Phi(\tilde{W}_{i})$, $Z_{i}=\Phi(\tilde{Z}_{i})$, $U_{i}=F_{U}^{-1}\left(\Phi(\tilde{U}_{i})\right)$, $(\tilde{W}_{i},\tilde{Z}_{i},\tilde{U}_{i})\sim N(0,\Sigma)$, $\Sigma_{11}=\Sigma_{22}=\Sigma_{33}=1$, $\Sigma_{23}=0$, and $\Phi$ is the standard normal CDF. Hence, in all the designs, both the regressors and the instruments are $\operatorname{Uniform}[0,1]$. We consider six DGPs that differ with respect to the error distribution $F_{U}$ and whether or not $W$ is exogenous.
@@
[…]
In Appendix F, we consider two additional DGPs to assess the impact of the strength of the instrument.
4.2 Bandwidth choice
An important practical issue is the choice of the bandwidths $h_{1,n}$, $h_{2,n}$, and $h_{3,n}$. We use the (data-dependent) rule-of-thumb bandwidth choice described in Section 3.4. To implement this rule-of-thumb bandwidth choice, we need to choose the constants $\tilde{A}_{j}$, $j\in\{G,Q,\kappa\}$. We found that across the DGPs considered in our simulations, $\tilde{A}_{G}=\tilde{A}_{\kappa}=2$ and $\tilde{A}_{Q}=1.5$ perform well in terms of bias across all designs, especially when the bias is large. [^8] If more information about the DGP is available, it is possible to further refine the tuning parameter. For example, under DGP1 where $Q=0$, choosing smaller values for $A_{Q}$ results in a better alignment with the infeasible formula. However, absent such information, we consider the proposed rule of thumb to be a reasonable compromise when one cares about the performance across different designs, especially those with larger biases (e.g., DGP3–DGP6).
We compare the performance of the feasible bias correction based on the rule-of-thumb bandwidth choice to the performance of the corresponding infeasible bias correction based on the true $G$, $Q$, and $\kappa$. For IVQR, the exact formulas are not available. Therefore, we use numerically computed values based on 10 million observations and tuning choice ${A}_{G}={A}_{Q}={A}_{\kappa}=1$.
4.3 Results
We focus on the performance of bias correction for $\tau\in\{0.25,0.5,0.75\}$. Figure 2 shows the impact of bias correction for the exogenous DGP1–DGP3 with $n=100$. Figure 3 shows the corresponding results for the endogenous DGPs (DGP4–DGP6). We use classical QR of $Y$ on $W$, implemented via the linear programming formulation in Appendix D, for DGP1–DGP3 and IVQR, implemented using the MILP formulation in Appendix D, for DGP4–DGP6.
The main findings can be summarized as follows. First, the QR estimators based on the exogenous designs (DGP1–DGP3) exhibit lower biases than the IVQR estimators based on the corresponding endogenous designs (DGP4–DGP6). For both estimators, the biases tend to be the largest for the designs with heavy tails (DGP3 and DGP6). Second, the feasible bias correction based on the rule-of-thumb bandwidth choice reduces the bias of QR and IVQR estimators in many cases. The bias reductions are the most notable when the biases of the original estimators are large, which is when bias correction is most needed. Finally, the infeasible bias correction reduces the bias in most cases, underscoring the usefulness of the proposed theory. See Appendix F for additional simulation evidence on the impact of the sample size and instrument strength on the performance of bias correction.
Next, we investigate the impact of bias correction on the RMSE of the estimators. The proposed bias correction approach is designed to reduce the bias but is not theoretically guaranteed to reduce the RMSE. Figure 4 reports the results for DGP1 and DGP4 with $n=100$. Overall, the impact of bias correction on the RMSE is rather small. While the infeasible bias correction can slightly increase the RMSE for DGP1, it decreases the RMSE across all quantile levels for DGP4. The feasible bias correction with the rule-of-thumb bandwidth slightly increases the RMSE in most cases. Appendix Figures 10 and 11 show the results for the other DGPs, and Appendix Figure 12 shows the corresponding results for the Mean Absolute Deviation, an alternative measure of risk.
Finally, we study the impact of the bias correction on the coverage probability of the standard confidence intervals. To focus on the impact of bias correction, we use the same standard errors based on the rule-of-thumb bandwidth for the original and the bias-corrected estimator. As a result, by construction, the bias correction does not affect the length of the confidence intervals. Figure 5 shows the empirical coverage for DGP1 and DGP4 before and after bias correction. The feasible bias correction based on the rule-of-thumb bandwidth leads to higher coverage accuracy in the majority of cases but can lead to some undercoverage at the median. Thus, while bias correction may not improve the RMSE in small samples, it can lead to more accurate inferences. The results for the other DGPs are in Appendix Figures 13 and 14, and Appendix Figure 15 presents the results for $n=200$.
Figure 2: Bias (multiplied by $n$) before and after correction for DGP1–DGP3